# Navigating in the Dark: A Multimodal Framework and Dataset for Nighttime Traffic Sign Recognition

### v1

@@ 4.1 LENS-Net Detector
1:Image crop $\mathbf{I}_{\text{crop}}$ (from LENS-Net Detector)
2:Frozen $E^{img}_{clip},\;E^{text}_{clip}$; Static text prompts $\mathbf{P}$ $\triangleright$ CLIP image and text encoders
3:Class logits $(\mathbf{I}_{\text{crop}})_{\text{class}}$
4:procedure ClassifierForward($\mathbf{I}_{\text{crop}}$) Get Foundational Embeddings
5:  $\mathbf{v}\leftarrow\text{L2Normalize}(E^{img}_{clip}(\mathbf{I}_{\text{crop}}))$ $\triangleright$ Image vector
6:  $\mathbf{T}\leftarrow\text{GetPrecomputedTextEmbeddings}(\mathbf{P})$ $\triangleright$ Prompt matrix Branch 1: Text-Image Fusion
7:  $\mathbf{f}_{\text{CA1}}\leftarrow\text{TextImageFusionModule}(\mathbf{v},\mathbf{T})$ $\triangleright$ Handles $\text{MHA}_{\text{CA1}}$, Res, LN, FFN (Equation 6) Branch 2: Graph Attribute Module
8:  $\mathbf{H}^{(L)}\leftarrow\text{GraphAttributeModule}(\mathbf{v})$ $\triangleright$ Handles node creation & GCNN Final Fusion & Classification
9:  $\mathbf{f}_{\text{final}}\leftarrow\text{FinalFusionModule}(\mathbf{f}_{\text{CA1}},\mathbf{H}^{(L)})$ $\triangleright$ Handles $\text{MHA}_{\text{CA2}}$, Res, LN, FFN (Equation 10)
10:  $(\mathbf{I}_{\text{crop}})_{\text{class}}\leftarrow\text{ClassifierHead}(\mathbf{f}_{\text{final}})$
11:  return $(\mathbf{I}_{crop})_{\text{class}}$ $\triangleright$ Output class category
12:end procedure
13:function GraphAttributeModule($\mathbf{v}$)
14:  $\mathbf{n}_{G}\leftarrow\mathbf{v}$
15:  $\mathbf{n}_{S}\leftarrow(\text{Softmax}(\text{MLP}_{S}(\mathbf{v})))\cdot\mathbf{E}_{S}$ $\triangleright$ $\mathbf{E}_{S}$: learnable shape embedding matrix
16:  $\mathbf{n}_{C}\leftarrow(\text{Softmax}(\text{MLP}_{C}(\mathbf{v})))\cdot\mathbf{E}_{C}$ $\triangleright$ $\mathbf{E}_{C}$: learnable color embedding matrix
17:  $\mathbf{H}^{(0)}\leftarrow\text{Stack}(\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C})$
18:  $\mathbf{H}^{(L)}\leftarrow\text{SimpleGCNN}(\mathbf{H}^{(0)})$
19:  return $\mathbf{H}^{(L)}$
20:end function
To ensure that preprocessing module learns meaningful transformations and avoids collapsing to trivial solutions, we employ two key strategies. (i) Stochastic Augmentation: During training, we apply a diverse set of strong, non-differentiable augmentations (photometric jitter, JPEG compression, Gaussian noise, blur, and small rotations) before the image is passed to the PreprocessWrapper. It forces $\Phi_{\text{head}}$ to learn to invert these degradations and normalize the image to a consistent domain for the detector. (ii) Preprocessing Loss: We apply an L2 regularization based loss ($\mathcal{L}_{\text{preproc}}$) that penalizes the predicted scaled parameters for deviating from their (no operation) default values.
$\mathcal{L}_{\text{preproc}}=\mathbb{E}\left[\sum_{i\in\{\gamma,\alpha,\zeta\}}\lambda_{i}(p_{i}-p_{\text{default},i})^{2}\right]$ (3)
where $\lambda_{i}$ ($\lambda_{i}\in(0,0.01)$) are per-parameter loss weights. This loss is added to the standard YOLOv8 detection loss, $\mathcal{L}_{\text{detect}}$, during training.
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{detect}}+\mathcal{L}_{\text{preproc}}$ (4)
LENS-Net detector is trained using a zero-one scheduler: we first freeze the backbone and train $\Phi_{\text{head}}$, then unfreeze the backbone for joint end-to-end training. The detector’s bounding-box output ($\mathbf{I}_{\text{crop}}$) is forwarded to the LENS-Net classifier, effectively filtering out the background, which does not contribute to TSR.
4.2 LENS-Net Classifier
After the detector localizes a traffic sign and outputs the corresponding cropped region, the classifier is tasked with accurately identifying the specific signboard category. We design a tri-modal fusion classifier that combines: (i) robust, global semantic features from a frozen CLIP vision encoder, (ii) prompt-conditioned textual priors from the CLIP text encoder, and (iii) structured attribute reasoning using a small graph. The complete architecture is detailed in Algorithm 1. The model processes the $\mathbf{I}_{\text{crop}}$ in two parallel branches that are fused for a final prediction.
Branch 1: Text-Image Fusion. We process $\mathbf{I}_{\text{crop}}$ to obtain the corresponding global image embedding $\mathbf{v}$, which is fused with semantic context derived from a set of pre-defined textual prompts $\mathbf{P}$. These prompts are encoded into embeddings $\mathbf{T}\in\mathbb{R}^{n(\mathbf{P})\times D}$ using the frozen CLIP text encoder (Algorithm 1, line 2-3). $\mathbf{v}$ serves as a holistic representation of the crop’s high-level visual semantics and $\mathbf{T}$ is used for every image processed by the classifier, acting as a ”semantic key/value store” that the model can attend to.
The L2 normalization of the image vector $\mathbf{v}$ and all $n(\mathbf{P})$ prompt vectors in $\mathbf{T}$ projects them onto the surface of a $D$-dimensional unit hypersphere, allowing their dot product to be directly interpreted as cosine similarity. Normalization also ensures that the magnitude does not dominate the later computations.
We use a cross-attention module, $\text{MHA}_{\text{CA1}}$ (Equation 5), where $\mathbf{v}$ acts as the query $(\mathbf{Q})$ and the text embeddings $\mathbf{T}$ act as keys $(\mathbf{K})$ and values $(\mathbf{V})$. This produces a prompt-conditioned image feature CA1, a refined global vector that already blends visual information with textual priors.
@@ 4.2 LENS-Net Classifier
The output $\mathbf{CA1}$ is processed through a residual connection (Res), Layer Normalization (LN), and position-wise feed-forward network (FFN) to get the text-fused image feature, $\mathbf{f}_{\text{CA1}}$, such that $\mathbf{f}_{\text{CA1}}\in\mathbb{R}^{D}$:
$\mathbf{f}_{\text{CA1}}=\text{FFN}(\text{LN}(\mathbf{v}+\mathbf{CA1}))$ (6)
Equation 5 and 6 represents the TextImageFusionModule in Algorithm 1 (line 4). $\mathbf{f}_{\text{CA1}}$ aligns the image’s representation with textual distinctions that are useful for traffic-sign recognition (e.g., color/shape descriptions, presence of numbers or arrows, etc.).
Branch 2: Graph Attribute Module. This branch explicitly models the two most salient attributes of signs: shape and color. As shown in the GraphAttributeModule (Algorithm 1, lines 10-17), we use the global image embedding $\mathbf{v}$ to predict soft distributions over learnable shape ($\mathbf{E}_{S}$) and color ($\mathbf{E}_{C}$) embedding tables, turning the predicted distributions into continuous node vectors, i.e., instead of making a hard, non-differentiable argmax decision (e.g., this is 100% a circle), we represent uncertainty (e.g., [0.85 (circle), 0.1 (octagon), 0.05 (other)]).
For fine-grained sign classification, we identify three pillars of information as critical: the global appearance, the geometric shape, and the dominant color. Using these pillars, we construct a 3-node graph $\mathbf{H}^{(0)}$, composed of global appearance ($\mathbf{n}_{G}=\mathbf{v}$), shape ($\mathbf{n}_{S}$), and color ($\mathbf{n}_{C}$) nodes. These three nodes (all designed to live in the same $D$-dimensional embedding space) form a structured 3-node graph: $\mathbf{H}^{(0)}=[\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C}]^{T}$. We refer the interested readers to Appendix 9 for more details on node construction.
We use a 3-layer GCNN (SimpleGCNN on line 15 in Algorithm 1) operating on node matrix $\mathbf{N}$ $(\mathbf{N}=[\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C}],\mathbf{N}\in\mathbb{R}^{D\times 3})$. Let $\mathbf{A}\in\mathbb{R}^{3\times 3}$ be the fully-connected adjacency matrix (every node talks to every other). We normalize the adjacency matrix symmetrically (Equation 7) and perform a linear mapping followed by propagation with $\hat{A}$ (Equation 8):
@@
where $\mathbf{H}^{(l)}\in\mathbb{R}^{3\times D}$ is node features at layer $l$, and $W^{(l)},b^{(l)}$ are the layer parameters. After $L$ layers, the GCNN outputs updated node features: $\mathbf{H}^{(L)}\in\mathbb{R}^{3\times D}.$
Figure 4: Fold-wise performance of LENS-Net for detection (left) and classification (right) on INTSD. A.P., A.R. and Acc. denotes macro-averaged precision, macro-averaged recall and accuracy, respectively.
The GCNN enables relational reasoning between nodes, i.e., if the $\mathbf{n}_{G}$ suggests an arrow while the $\mathbf{n}_{S}$ indicates an octagon, the GCNN mediates these signals to produce a joint representation capturing their agreement or conflict. Since the $\mathbf{n}_{G}$ and $\mathbf{n}_{S}$ originate from shared global embeddings, the GCNN effectively refines their features through message passing.
Final Fusion. After the GCNN, we fuse the two branches using a second cross-attention module, $\text{MHA}_{\text{CA2}}$ (Algorithm 1, line 6). Here, the text-fused feature $\mathbf{f}_{\text{CA1}}$ queries the GCNN-refined attribute nodes $\mathbf{H}^{(L)}$ (Equation 9). This is a deliberate design choice: CA1 lets textual priors refine the image and CA2 lets the refined image go back and attend to structured attribute features. Conceptually, CA2 asks: “given what I now know from text, and given the shape/color-conditioned node features, how should I pool the node information to make final decision?” Mathematically,
@@
This output is again processed with a residual connection, LN, and an FFN to produce the final, fully-fused feature vector $\mathbf{f}_{\text{final}}$:
@@
The final vector $\mathbf{f}_{\text{final}}$ (Equation 10) is passed to an MLP head (Algorithm 1, line 7) and produces logits ($(\mathbf{I}_{\text{crop}})_{\text{class}}\in\mathbb{R}^{N}$) over the target traffic-sign classes.
Classifier Training Objective. The LENS-Net classifier is trained using the focal loss [25] that includes per-class weights computed from training frequencies (rarer classes get larger weights to reduce bias toward common classes). This is done to address the significant class imbalance inherent in INTSD.
@@
where $p_{t}$ is the model’s predicted probability for the ground-truth class, and $\gamma$ is the focusing parameter. We further employ a class-balancing variant where $\alpha_{t}$ is the inverse square root of the class frequency, normalized to sum to $N$.
$\alpha_{t}=\frac{1/\sqrt{\text{count}(t)}}{\sum_{j=1}^{N}(1/\sqrt{\text{count}(j)})}\times N$ (12)
To further mitigate imbalance, we employ a rarity-based augmentation (mentioned in Appendix 9). Implementation details are given in Appendix 8.
@@ 5 Results
Experimental Settings. We adopt a startified 5-fold cross-validation strategy for training and evaluation. We choose mAP@50 and mAP@50:95 as detection evaluation metrics and macro-averaged precision (A.P.), macro-averaged recall (A.R.), and accuracy (Acc.) as classification evaluation metrics.
Baseline Results. In Tables 2 and 3, we show strong baselines and benchmark INTSD for detection and classification, respectively. We demonstrate experimental results using: (i) DAI-Net [7], (ii) Faster R-CNN [39] with Feature Pyramid Network (FPN) [21] and residual networks (ResNet) [17] as their backbone, (iii) YOLOv8 [37], (iv) YOLO-TS [3], (v) BLIP [23], (v) CLIP [36], (vi) DeIT [47] and Rishabh et al. [50].
@@
Detection. Table 2 summarizes the detection performance on the INTSD dataset. The proposed LENS-Net detector achieves the highest accuracy among all compared methods, substantially outperforming Faster R-CNN [39] and DAI-Net [7]. While YOLOv8 [37] and YOLO-TS [3] show comparable results, both lag behind LENS-Net, confirming the benefit of our adaptive enhancement and the effectiveness of our detector in handling diverse and challenging traffic sign instances.
Figure 4 (left) shows the 5-fold performance of the LENS-Net detector. The scores are tightly clustered around 92.5% mAP@50, indicating strong and consistent detection ability. The smaller mAP@50:95 values suggest slightly lower precision in tight localization, though the low variance demonstrates overall stability.
Table 2: Detection performance on the INTSD.
Model mAP@50 (%) mAP@50:95 (%)
Faster R-CNN [39] 62.60 43.22
DAI-Net [7] 87.18 62.56
YOLOv8 [37] 90.13 77.21
YOLO-TS [3] 90.77 77.27
LENS-Net 92.56 79.54
Classification. Table 3 presents the classification results on INTSD. LENS-Net achieves the best performance across all metrics, outperforming the previous state-of-the-art by Rishabh et al. [50] by +8.4% in A.P. and +8.57% in A.R. It also yields the highest accuracy of 88.59%, highlighting the strength of its multimodal fusion and attribute-guided reasoning.
LENS-Net classifier shows consistently high accuracy (Figure 4, right), with scores tightly clustered around 88.5%, indicating stability across folds. In contrast, A.P. and A.R. show lower medians ($\approx$ 79.5% and 75.5%) and higher variance, suggesting less consistency in handling per-class precision-recall performance. Figure 5 gives the predictions of LENS-Net on INTSD along with false positives and false negatives.
Table 3: Classification results on the INTSD.
Model A.P. (%) A.R. (%) Acc. (%)
BLIP [23] 49.33 47.81 69.90
CLIP [36] 65.25 68.11 82.49
DeIT [47] 65.89 59.69 82.13
Rishabh et al. [50] 70.49 66.99 85.21
LENS-Net 78.89 75.56 88.59
Ablation Study. Figure 6 depicts the contribution of key components in LENS-Net. Removing the GraphAttributeModule causes a 3.92% and 3.36% drop in A.P. and A.R., confirming that the GCNN effectively models relational dependencies among global, shape, and color nodes through message passing, enabling compositional reasoning (e.g., “red octagon $\rightarrow$ stop sign”).
Replacing cross-attention fusion (TextImageFusionModule and FinalFusionModule) with simple concatenation reduces A.P. and A.R. by 5.0% and 3.18%, validating that cross-attention is crucial for meaningful multimodal interaction among image, text, and graph features. Similarly, ablating text prompts ($\mathbf{P}$) leads to a 2.21% and 1.84% drop, showing their role as text-guided priors that inject semantic knowledge of sign attributes and leverage CLIP’s vision-language alignment to resolve visually similar classes.
Eliminating the learned shape ($\mathbf{E}_{S}$) and color ($\mathbf{E}_{C}$) embedding tables and using simple projections ($MLP_{S}(\mathbf{v})$ and $MLP_{C}(\mathbf{v})$) results in 3.8% and 1.86% drops in A.P. and A.R. These embeddings allow the model to learn semantically rich geometric and chromatic representations within the CLIP feature space.
Figure 6: Ablation study on the impact of major components in LENS-Net for traffic sign classification on INTSD.

### v2

@@ 4.2 LENS-Net Classifier
1:Image crop $\mathbf{I}_{\text{crop}}$ (from LENS-Net Detector) 2:Frozen $E^{img}_{clip},\;E^{text}_{clip}$; Static text prompts $\mathbf{P}$ $\triangleright$ CLIP encoders 3:Class logits $(\mathbf{I}_{\text{crop}})_{\text{class}}$ 4:procedure ClassifierForward($\mathbf{I}_{\text{crop}}$) Get Foundational Embeddings 5:  $\mathbf{v}\leftarrow\text{L2Normalize}(E^{img}_{clip}(\mathbf{I}_{\text{crop}}))$ $\triangleright$ Image vector 6:  $\mathbf{T}\leftarrow\text{GetPrecomputedTextEmbeddings}(\mathbf{P})$ $\triangleright$ Prompt matrix Branch 1: Text-Image Fusion 7:  $\mathbf{f}_{\text{CA1}}\leftarrow\text{TextImageFusionModule}(\mathbf{v},\mathbf{T})$ $\triangleright$ Equation 2 Branch 2: Graph Attribute Module 8:  $\mathbf{H}^{(L)}\leftarrow\text{GraphAttributeModule}(\mathbf{v})$ $\triangleright$ Handles node creation & GCNN Final Fusion & Classification 9:  $\mathbf{f}_{\text{final}}\leftarrow\text{FinalFusionModule}(\mathbf{f}_{\text{CA1}},\mathbf{H}^{(L)})$ $\triangleright$ Equation 6 10:  $(\mathbf{I}_{\text{crop}})_{\text{class}}\leftarrow\text{ClassifierHead}(\mathbf{f}_{\text{final}})$ 11:  return $(\mathbf{I}_{crop})_{\text{class}}$ $\triangleright$ Output class category 12:end procedure 13:function GraphAttributeModule($\mathbf{v}$) 14:  $\mathbf{n}_{G}\leftarrow\mathbf{v}$ 15:  $\mathbf{n}_{S}\leftarrow(\text{Softmax}(\text{MLP}_{S}(\mathbf{v})))\cdot\mathbf{E}_{S}$ $\triangleright$ $\mathbf{E}_{S}$: learnable shape embedding matrix 16:  $\mathbf{n}_{C}\leftarrow(\text{Softmax}(\text{MLP}_{C}(\mathbf{v})))\cdot\mathbf{E}_{C}$ $\triangleright$ $\mathbf{E}_{C}$: learnable color embedding matrix 17:  $\mathbf{H}^{(0)}\leftarrow\text{Stack}(\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C})$ $\triangleright$ Graph construction 18:  $\mathbf{H}^{(L)}\leftarrow\text{SimpleGCNN}(\mathbf{H}^{(0)})$ $\triangleright$ Updated node features 19:  return $\mathbf{H}^{(L)}$ 20:end function
@@
$\mathbf{f}_{\text{CA1}}=\text{FFN}(\text{LN}(\mathbf{v}+\mathbf{CA1})),\>\>\mathbf{f}_{\text{CA1}}\in\mathbb{R}^{D}$ (2)
Equation 1 and 2 represents the TextImageFusionModule in Algorithm 1 (line 4). $\mathbf{f}_{\text{CA1}}$ aligns the image’s representation with textual distinctions that are useful for traffic-sign recognition (e.g., color/shape descriptions, presence of numbers or arrows, etc.).
Branch 2: Graph Attribute Module. This branch explicitly models the two most salient attributes of signs: shape and color. As shown in the GraphAttributeModule (Algorithm 1, lines 10-17), from the global embedding $\mathbf{v}$, we predict soft distributions over learnable shape ($\mathbf{E}_{S}$) and color ($\mathbf{E}_{C}$) embedding tables. It turns the predicted distributions into continuous node vectors, i.e., instead of making a hard, non-differentiable argmax decision (e.g., this is 100% a circle), we represent uncertainty (e.g., [0.85 (circle), 0.1 (octagon), 0.05 (other)]).
For fine-grained sign classification, we construct three nodes: global ($\mathbf{n}_{G}$), shape ($\mathbf{n}_{S}$), and color ($\mathbf{n}_{C}$). Additional details on node construction are provided in Section 3 of the supplementary material. These form a structured 3-node graph: $\mathbf{H}^{(0)}=[\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C}]^{T}$, refined using a 3-layer GCNN (SimpleGCNN on line 15 in Algorithm 1) operating on node matrix $\mathbf{N}$ $(\mathbf{N}=[\mathbf{n}_{G},\mathbf{n}_{S},\mathbf{n}_{C}],\mathbf{N}\in\mathbb{R}^{D\times 3})$. Let $\mathbf{A}\in\mathbb{R}^{3\times 3}$ be the fully-connected and symmetrically normalized adjacency (Equation 34). The GCNN outputs updated node features $\mathbf{H}^{\text{L}}$.
@@
where $\mathbf{H}^{(l)}\in\mathbb{R}^{3\times D}$ is node features at layer $l$, and $W^{(l)},b^{(l)}$ are the layer parameters. After $L$ layers, the GCNN outputs node features: $\mathbf{H}^{(L)}\in\mathbb{R}^{3\times D}.$
Final Fusion. The text-conditioned feature $\mathbf{f}_{\text{CA1}}$ queries the GCNN-refined node features $\mathbf{H}^{(L)}$ via a second cross-attention module (Equation 5; Algorithm 1, line 6). CA1 lets textual priors refine the image and CA2 lets the refined image go back and attend to structured attribute features. The output is processed with residual normalization and FFN to obtain the final feature $\mathbf{f}_{\text{final}}$ (Equation 6). The final vector $\mathbf{f}_{\text{final}}$ is passed to an MLP head (Algorithm 1, line 7) and produces logits ($(\mathbf{I}_{\text{crop}})_{\text{class}}\in\mathbb{R}^{N}$) over the target traffic-sign classes.
@@
Classifier Training Objective. The classifier is trained using class-balanced focal loss [25] (Equation 7) to mitigate long-tailed class imbalance in INTSD. Class weights $\alpha_{t}$ are set proportional to the inverse square root of class frequency.
@@
where $p_{t}$ is the model’s predicted probability for the ground-truth class, and $\gamma$ is the focusing parameter. To further mitigate imbalance, we employ a rarity-based augmentation. Implementation details and augmentation strategies are mentioned in the supplementary material (Section 2).
@@ 5 Results
Experimental Settings. We adopt a stratified 5-fold cross-validation strategy for training and evaluation, thereby reducing bias in the evaluation process. We choose mAP@50 and mAP@50:95 as detection evaluation metrics and macro-averaged precision (A.P.), macro-averaged recall (A.R.), and accuracy (Acc.) as classification evaluation metrics. We also report frames per second (FPS) to evaluate real-time performance and deployment feasibility. For fair comparison, FPS is measured end-to-end, from raw image loading to final predictions, including preprocessing, forward pass, post-processing, and CUDA synchronization. Each FPS value is averaged over five independent runs.
Baseline Results. In Tables 2 and 3, we benchmark INTSD using a diverse set of strong baselines covering multiple architectural paradigms. For detection, we evaluate: (i) two-stage detectors (Faster R-CNN [40]), (ii) one-stage detectors (YOLOv8 [18], YOLOv10-L [53], YOLOv12-L [48], GOLD-YOLO-L [54]), (iii) low-light–oriented detection methods (DAI-Net [7], YOLO-LLTS [27], YOLO-TS [3]), and (iv) TSR approaches (YOLO-LLTS [27], YOLO-TS [3]).
For classification, we compare against: (i) convolutional and transformer-based models (DeIT [49]), (ii) vision-language models (CLIP [37], BLIP [23]), (iii) YOLO based methods (YOLOv10-L [53], YOLOv12-L [48]) and (iv) task-specific TSR approaches (Rishabh et al. [52], YOLO-TS [3]). All models are trained and evaluated under identical data splits and reporting protocols.
Detection. Table 2 reports detection performance on INTSD across two-stage, one-stage, and low-light–oriented detectors. Modern YOLO variants achieve strong results, reflecting the dataset’s suitability for contemporary detection architectures. LENS-Net attains the highest mAP@50 (92.56%) and mAP@50:95 (79.54%), while maintaining competitive inference speed. The consistent performance gap over both generic and low-light–specific detectors highlights the challenges posed by diverse nighttime conditions and validates the benchmark’s ability to differentiate strong methods.
Table 2: Detection performance on the INTSD.
Model mAP@50 (%) mAP@50:95 (%) FPS Parameters (M)
Total Trainable
Faster R-CNN [40] 62.60 43.22 23.38 41.50 41.50
DAI-Net [7] 87.18 62.56 20.15 49.83 49.83
YOLOv12-L [48] 82.84 64.93 59.08 26.48 26.48
GOLD-YOLO-L [54] 88.70 72.11 42.64 74.99 74.99
YOLO-LLTS [27] 90.00 73.65 56.34 9.87 9.87
YOLOv10-L [53] 90.00 75.81 64.39 25.83 25.83
YOLOv8 [18] 90.13 77.21 83.71 11.14 11.14
YOLO-TS [3] 90.82 77.27 62.52 23.39 23.39
LENS-Net 92.56 79.54 82.27 11.22 11.22
Table 3: Classification results on the INTSD.
Model A.P. (%) A.R. (%) Acc. (%) FPS Parameters (M)
Total Trainable
YOLOv10-L [53] 37.24 26.51 70.32 64.37 25.83 25.83
YOLOv12-L [48] 45.70 39.60 67.66 59.14 26.48 26.48
BLIP [23] 49.33 47.81 69.90 75.37 4060.87 8.93
CLIP [37] 65.25 68.11 82.49 72.37 434.04 12.85
DeIT [49] 65.89 59.69 82.13 136.98 85.83 85.83
Rishabh et al. [52] 74.49 69.98 88.21 60.25 26.87 26.85
YOLO-TS [3] 75.08 74.43 89.70 63.36 23.39 23.39
LENS-Net 78.89 75.56 88.59 62.44 439.37 11.75
Classification. Table 3 presents classification performance across convolutional, transformer-based, vision-language, and task-specific TSR models. Vision-language approaches such as CLIP and BLIP provide strong semantic baselines, while task-specific models (e.g., Rishabh et al. [52], YOLO-TS [3]) achieve higher task-adapted accuracy. LENS-Net achieves the best macro-averaged precision (78.89%) and recall (75.56%), demonstrating the effectiveness of structured multimodal reasoning under long-tailed and low-illumination conditions. Ablation study (see supplementary material, Section 4) confirms the contribution of key components in LENS-Net classifier. These results establish competitive reference baselines for future research on INTSD. Figure 5 gives the predictions of LENS-Net on INTSD along with false positives and false negatives.
@@
Generalization to CNTSSS. Table 4 reports cross-dataset evaluation on CNTSSS [27]. While several methods achieve strong performance due to its limited class taxonomy, INTSD produces larger performance variation across models, reflecting increased semantic diversity and intra-class ambiguity. Notably, LENS-Net achieves competitive detection performance and the highest classification accuracy, demonstrating robustness across differing nighttime benchmarks.
Table 4: Detection and classification performance on CNTSSS. Rishabh et al. and YOLO-LLTS are classification-only and detection-only models, respectively.
Model Detection (%) Classification (%)
mAP@50 mAP@50:95 FPS Parameters (M) A.P. A.R. Acc. FPS Parameters (M)
YOLOv10-L 68.84 46.56 100.51 26.77 84.19 84.15 86.02 100.32 26.77
Rishabh et al. 90.00 89.33 91.68 63.52 26.85
YOLO-LLTS 76.30 54.15 72.17 9.85
YOLO-TS 76.47 54.21 102.51 23.39 83.92 84.08 86.46 102.54 23.39
LENS-Net 69.17 46.62 151.75 11.13 90.67 90.02 92.10 70.60 11.73